vector R = vector P + vector Q
if P = Q = R, find the angle between P and Q.

Grade:11

1 Answers

Abhishek singh
6 years ago
Let P=Q=R=AMagnitude of vectorR is|R|=(P^2+Q^2+2PQcosthita)where thita is angle between p and qA^2=A^2+A^2+2A^2costhitaCosthita=-1/2Thita=120*Hence angle between p and q is 120 degree